# Heavy Mesons to Charmed Tetraquark Decays

### v1

@@ IV.1.1 Topological amplitudes in $\overline{B}\to D\overline{T}$ decays
and with $T_{\bar{T}},C_{\bar{T}},E_{\bar{T}\bar{B}}$ replaced with $T^{\prime}_{\bar{T}},C^{\prime}_{\bar{T}},E^{\prime}_{\bar{T}\bar{B}}$, respectively. It is understood that these amplitudes are related by Cabibbo–Kobayashi–Maskawa (CKM) matrix elements, giving $T^{\prime}_{\bar{T}}=\frac{V^{*}_{cd}}{V^{*}_{cs}}T_{\bar{T}}$, $C^{\prime}_{\bar{T}}=\frac{V^{*}_{cd}}{V^{*}_{cs}}C_{\bar{T}}$ and $E^{\prime}_{\bar{T}\bar{B}}=\frac{V^{*}_{cd}}{V^{*}_{cs}}E_{\bar{T}\bar{B}}$.
Figure 5: Topological diagrams of (a) $C_{\bar{D}T}$ (internal $W$), (b) $T_{\bar{D}}$ (external $W$) and (c) $E_{\bar{D}\bar{B}}$ (exchange), amplitudes in $\overline{B}\to\overline{D}T$ decays. The subscripts indicate the final state meson(s) receiving the $\bar{c}q_{j}$ quarks from the $W$ line in the $b\to cW^{-},W^{-}\to\bar{c}q_{j}$ transition. These are flavor flow diagrams. Only (a) contributes to modes with flavor exotic states or other states in the same multiplets.
@@ IV.1.2 Topological amplitudes in $\overline{B}\to\overline{D}T$ decays
We turn to $\overline{B}_{q}\to\overline{D}T$ decays. The $b\to cW^{-},W^{-}\to\bar{c}s$ transition can also induce $\overline{B}\to\overline{D}T(cq[\bar{q}^{\prime}\bar{q}^{\prime\prime}])$ and $\overline{B}\to\overline{D}T(cq\{\bar{q}^{\prime}\bar{q}^{\prime\prime}\})$ decays as well. The Hamiltonian governing $\overline{B}\to\overline{D}T(cq[\bar{q}^{\prime}\bar{q}^{\prime\prime}])$ decays in $\Delta S=-1$ transitions in scenario I is given by
$H_{\rm eff}$ $=$ $T_{\bar{D}}\,\overline{B}_{m}\overline{T}_{i}^{[mi]}H^{j}D_{j}+C_{\bar{D}T}\,\overline{B}_{m}H^{j}\overline{T}_{j}^{[mk]}D_{k}+E_{\bar{D}\bar{B}}\,\overline{B}_{i}H^{i}\overline{T}_{j}^{[jk]}D_{k},$ (38)
while the one for $\overline{B}\to D\overline{T}(\bar{c}\bar{q}\{q^{\prime}q^{\prime\prime}\})$ decays in scenario II is given by,
$H_{\rm eff}$ $=$ $T_{\bar{D}}\,\overline{B}_{m}\overline{T}_{i}^{\{mi\}}H^{j}D_{j}+C_{\bar{D}T}\,\overline{B}_{m}H^{j}\overline{T}_{j}^{\{mk\}}D_{k}+E_{\bar{D}\bar{B}}\,\overline{B}_{i}H^{i}\overline{T}_{j}^{\{jk\}}D_{k}.$ (39)
@@
For notational simplicity, we use the same notations on topological amplitudes in both scenarios. There are three topological amplitudes, $T_{\bar{D}}$, $C_{\bar{D}T}$ and $E_{\bar{D}\bar{B}}$ and they are depicted in Fig. 5. These $T_{\bar{D}}$, $C_{\bar{D}T}$ and $E_{\bar{D}\bar{B}}$ amplitudes denote external $W$, internal $W$ and exchange amplitudes, respectively, and the subscripts denote the particle(s) receiving the $\bar{c}s$ quarks from the $W$ line in the $b\to cW^{-},W^{-}\to\bar{c}s$ transition. The effective Hamiltonians governing $\Delta S=0$ transitions can be obtained from the above equations with $H$ replaced with $H^{\prime}$ given in Eq. (37), and with $T_{\bar{D}},C_{\bar{D}T},E_{\bar{D}\bar{B}}$ replaced with $T^{\prime}_{\bar{D}},C^{\prime}_{\bar{D}},E^{\prime}_{\bar{D}\bar{B}}$, respectively. It is understood that these amplitudes are related by CKM matrix elements, giving $T^{\prime}_{\bar{D}}=\frac{V^{*}_{cd}}{V^{*}_{cs}}T_{\bar{D}}$, $C^{\prime}_{\bar{D}T}=\frac{V^{*}_{cd}}{V^{*}_{cs}}C_{\bar{D}T}$ and $E^{\prime}_{\bar{D}\bar{B}}=\frac{V^{*}_{cd}}{V^{*}_{cs}}E_{\bar{D}\bar{B}}$.
The relative sizes of these topological amplitudes are expected to be
$|T_{\bar{D}}|>|C_{\bar{D}T}|>|E_{\bar{D}\bar{B}}|.$ (40)
@@
Although the external $W$ diagram is expected to dominate, it is unable to produce a flavor exotic $T$, as the $q\bar{q}$ pair in the diagram, see Fig. 5 (b), is flavor connected. The $\overline{B}_{q}\to\overline{D}T$ decay amplitudes of modes involving $T$ in ${\bf 6}$ in scenario I or $\overline{\bf 15}$ in scenario II only have contribution from $C_{\bar{D}T}$, i.e. the diagram depicted in 5 (a). Consequently, the amplitudes of these states are highly related.
@@
From the above discussion, we may wrongly have the impression that a $\overline{B}\to\overline{D}T$ decay is similar to a $\overline{B}\to D\overline{T}$ decay. In fact, by comparing Fig. 5 to Fig. 4, one can easily see the differences. For example, in $\overline{B}\to D\overline{T}$ decays, the external $W$ diagram, $T_{\bar{T}}$, has a $\overline{T}$ emitted from $W^{-}$, as shown in Fig. 4 (b), while in $\overline{B}\to\overline{D}T$ decays, the emitted particle is $\overline{D}$ instead, see Fig. 5 (b). Furthermore, there is no counterpart of the internal $W$ diagram $C_{\bar{D}T}$, Fig. 5 (a), in $\overline{B}\to D\overline{T}$ decays, as the $C_{\bar{T}}$ in Fig. 4 (a) is different. Explicitly, the former has $\bar{D}T$ sharing the $c\bar{q}$ from $W$, while in the latter, both $c\bar{q}$ go to $\bar{T}$. Consequently, one cannot identify the topological amplitudes in $\overline{B}\to D\overline{T}$ to those in $\overline{B}\to\overline{D}T$ decays.
@@ IV.2.1 $\overline{B}\to D\overline{T}$ decay amplitudes
# Mode $A(\overline{B}_{q}\to D\overline{T}_{\bar{c}\bar{q}[q^{\prime}q^{\prime\prime}]})$ Mode $A(\overline{B}_{q}\to D\overline{T}_{\bar{c}\bar{q}[q^{\prime}q^{\prime\prime}]})$
$\bar{1}^{*}$ $B^{-}\to D^{+}\overline{T}^{\,--}_{\bar{c}s}$ $-C_{\bar{T}}$ $\bar{2}^{*}$ $B^{-}\to D^{0}\overline{T}^{\,-}_{\bar{c}s}$ $-\frac{1}{\sqrt{2}}C_{\bar{T}}$
$\bar{2}^{*}$ $\overline{B}{}^{0}\to D^{+}\overline{T}^{\,-}_{\bar{c}s}$ $\frac{1}{\sqrt{2}}C_{\bar{T}}$ $\bar{3}^{*}$ $\overline{B}{}^{0}\to D^{0}\overline{T}^{0}_{\bar{c}s}$ $-C_{\bar{T}}$
$\bar{5}$ $\overline{B}{}_{s}^{0}\to D^{+}\overline{T}^{\,-}_{\bar{c}}$ $-\frac{1}{\sqrt{2}}C_{\bar{T}}$ $\bar{6}$ $\overline{B}{}_{s}^{0}\to D^{0}\overline{T}^{0}_{\bar{c}}$ $\frac{1}{\sqrt{2}}C_{\bar{T}}$
$\bar{7}$ $\overline{B}{}_{s}^{0}\to D^{+}\overline{T}^{\,\prime\prime-}_{\bar{c}}$ $\frac{1}{\sqrt{2}}(C_{\bar{T}}+2E_{\bar{T}\bar{B}})$ $\bar{8}$ $\overline{B}{}_{s}^{0}\to D^{0}\overline{T}^{\,\prime\prime 0}_{\bar{c}}$ $\frac{1}{\sqrt{2}}(C_{\bar{T}}+2E_{\bar{T}\bar{B}})$
$\bar{9}$ $B^{-}\to D^{0}\overline{T}^{\,\prime\prime-}_{\bar{c}s}$ $-\frac{1}{\sqrt{2}}(C_{\bar{T}}+2T_{\bar{T}})$ $\bar{9}$ $\overline{B}{}^{0}\to D^{+}\overline{T}^{\,\prime\prime-}_{\bar{c}s}$ $-\frac{1}{\sqrt{2}}(C_{\bar{T}}+2T_{\bar{T}})$
$\bar{9}$ $\overline{B}{}_{s}^{0}\to D_{s}^{+}\overline{T}^{\,\prime\prime-}_{\bar{c}s}$ $\sqrt{2}(-T_{\bar{T}}+E_{\bar{T}\bar{B}})$
# Mode $A^{\prime}(\overline{B}_{q}\to D\overline{T}_{\bar{c}\bar{q}[q^{\prime}q^{\prime\prime}]})$ Mode $A^{\prime}(\overline{B}_{q}\to D\overline{T}_{\bar{c}\bar{q}\{q^{\prime}q^{\prime\prime}\}})$
$\bar{1}^{*}$ $B^{-}\to D_{s}^{+}\overline{T}^{\,--}_{\bar{c}s}$ $C^{\prime}_{\bar{T}}$ $\bar{2}^{*}$ $\overline{B}{}^{0}\to D_{s}^{+}\overline{T}^{\,-}_{\bar{c}s}$ $-\frac{C^{\prime}_{\bar{T}}}{\sqrt{2}}$
$\bar{4}^{*}$ $\overline{B}{}_{s}^{0}\to D^{0}\overline{T}^{0}_{\bar{c}\bar{s}}$ $-C^{\prime}_{\bar{T}}$
$\bar{5}$ $B^{-}\to D^{0}\overline{T}^{\,-}_{\bar{c}}$ $-\frac{1}{\sqrt{2}}C^{\prime}_{\bar{T}}$ $\bar{5}$ $\overline{B}{}_{s}^{0}\to D_{s}^{+}\overline{T}^{\,-1}_{\bar{c}}$ $\frac{1}{\sqrt{2}}C^{\prime}_{\bar{T}}$
$\bar{6}$ $\overline{B}{}^{0}\to D^{0}\overline{T}^{0}_{\bar{c}}$ $-\frac{1}{\sqrt{2}}C^{\prime}_{\bar{T}}$
$\bar{7}$ $B^{-}\to D^{0}\overline{T}^{\,\prime\prime-}_{\bar{c}}$ $-\frac{1}{\sqrt{2}}(C^{\prime}_{\bar{T}}+2T^{\prime}_{\bar{T}})$ $\bar{7}$ $\overline{B}{}^{0}\to D^{+}\overline{T}^{\,\prime\prime-}_{\bar{c}}$ $\sqrt{2}(E^{\prime}_{\bar{T}\bar{B}}-T^{\prime}_{\bar{T}})$
$\bar{7}$ $\overline{B}{}_{s}^{0}\to D_{s}^{+}\overline{T}^{\,\prime\prime-}_{\bar{c}}$ $-\frac{1}{\sqrt{2}}(C^{\prime}_{\bar{T}}+2T^{\prime}_{\bar{T}})$ $\bar{8}$ $\overline{B}{}^{0}\to D^{0}\overline{T}^{\,\prime\prime 0}_{\bar{c}}$ $\frac{1}{\sqrt{2}}(C^{\prime}_{\bar{T}}+2E^{\prime}_{\bar{T}\bar{B}})$
$\bar{9}$ $\overline{B}{}^{0}\to D_{s}^{+}\overline{T}^{\,\prime\prime-}_{\bar{c}s}$ $\frac{1}{\sqrt{2}}(C^{\prime}_{\bar{T}}+2E^{\prime}_{\bar{T}\bar{B}})$

### v3

@@ IV.1.1 Topological amplitudes in $\overline{B}\to D\overline{T}$ decays
and with $T_{\bar{T}}$, $C_{\bar{T}}$, $E_{\bar{T}\bar{B}}$, $P_{\bar{T},1}$, $P_{\bar{T},2}$, $PA_{\bar{B}}$ replaced with $T^{\prime}_{\bar{T}}$, $C^{\prime}_{\bar{T}}$, $E^{\prime}_{\bar{T}\bar{B}}$, $P^{\prime}_{\bar{T},1}$, $P^{\prime}_{\bar{T},2}$, $PA^{\prime}_{\bar{B}}$ respectively. It is understood that these amplitudes are related by Cabibbo–Kobayashi–Maskawa (CKM) matrix elements, giving $T^{\prime}_{\bar{T}}=\frac{V^{*}_{cd}}{V^{*}_{cs}}T_{\bar{T}}$, $C^{\prime}_{\bar{T}}=\frac{V^{*}_{cd}}{V^{*}_{cs}}C_{\bar{T}}$, $E^{\prime}_{\bar{T}\bar{B}}=\frac{V^{*}_{cd}}{V^{*}_{cs}}E_{\bar{T}\bar{B}}$, $P^{\prime}_{\bar{T},1}=\frac{V^{*}_{td}}{V^{*}_{ts}}P_{\bar{T},1}$, $P^{\prime}_{\bar{T},2}=\frac{V^{*}_{td}}{V^{*}_{ts}}P_{\bar{T},2}$ and $PA^{\prime}_{\bar{B}}=\frac{V^{*}_{td}}{V^{*}_{ts}}PA_{\bar{B}}$.
Figure 5: Topological diagrams of (a) $C_{\bar{D}T}$ (internal $W$), (b) $T_{\bar{D}}$ (external $W$), (c) $E_{\bar{D}\bar{B}}$ (exchange), (d) $P_{T}$ (penguin), (e) $P_{\bar{D}}$ (penguin) and (f) $PA_{\bar{B}}$ (penguin annihilation) amplitudes in $\overline{B}\to\overline{D}T$ decays. These are flavor flow diagrams. Only (a) and (d) contribute to modes with flavor exotic states or other states in the same multiplets.
@@ IV.1.2 Topological amplitudes in $\overline{B}\to\overline{D}T$ decays
We turn to $\overline{B}_{q}\to\overline{D}T$ decays. The tree $b\to cW^{-},W^{-}\to\bar{c}s$ and the penguin $b\to s\bar{c}c$ transitions can also induce $\overline{B}\to\overline{D}T(cq[\bar{q}^{\prime}\bar{q}^{\prime\prime}])$ and $\overline{B}\to\overline{D}T(cq\{\bar{q}^{\prime}\bar{q}^{\prime\prime}\})$ decays as well. The Hamiltonian governing the $\overline{B}\to\overline{D}T(cq[\bar{q}^{\prime}\bar{q}^{\prime\prime}])$ decays in $\Delta S=-1$ transitions in scenario I is given by
$H_{\rm eff}$ $=$ $T_{\bar{D}}\,\overline{B}_{m}\overline{T}_{i}^{[mi]}H^{j}D_{j}+C_{\bar{D}T}\,\overline{B}_{m}H^{j}\overline{T}_{j}^{[mk]}D_{k}+E_{\bar{D}\bar{B}}\,\overline{B}_{i}H^{i}\overline{T}_{j}^{[jk]}D_{k}$ (54)
$+P_{T}\,\overline{B}_{m}H^{j}\overline{T}_{j}^{[mk]}D_{k}+P_{\bar{D}}\,\overline{B}_{m}\overline{T}_{i}^{[mi]}H^{j}D_{j}+PA_{\bar{B}}\,\overline{B}_{i}H^{i}\overline{T}_{j}^{[jk]}D_{k},$
while the one for the $\overline{B}\to D\overline{T}(\bar{c}\bar{q}\{q^{\prime}q^{\prime\prime}\})$ decays in scenario II is given by,
$H_{\rm eff}$ $=$ $T_{\bar{D}}\,\overline{B}_{m}\overline{T}_{i}^{\{mi\}}H^{j}D_{j}+C_{\bar{D}T}\,\overline{B}_{m}H^{j}\overline{T}_{j}^{\{mk\}}D_{k}+E_{\bar{D}\bar{B}}\,\overline{B}_{i}H^{i}\overline{T}_{j}^{\{jk\}}D_{k}$ (55)
$+P_{T}\,\overline{B}_{m}H^{j}\overline{T}_{j}^{\{mk\}}D_{k}+P_{\bar{D}}\,\overline{B}_{m}\overline{T}_{i}^{\{mi\}}H^{j}D_{j}+PA_{\bar{B}}\,\overline{B}_{i}H^{i}\overline{T}_{j}^{\{jk\}}D_{k}.$
@@
For notational simplicity, we use the same notations on topological amplitudes in both scenarios. There are three tree topological amplitudes, $T_{\bar{D}}$, $C_{\bar{D}T}$ and $E_{\bar{D}\bar{B}}$ and three penguin topological amplitudes, $P_{T}$, $P_{\bar{D}}$ and $PA_{\bar{B}}$, and they are depicted in Fig. 5. These $T_{\bar{D}}$, $C_{\bar{D}T}$ and $E_{\bar{D}\bar{B}}$ amplitudes denote external $W$, internal $W$ and exchange amplitudes, respectively, and the subscripts denote the particle(s) receiving the $\bar{c}s$ quarks from the $W$ line in the $b\to cW^{-},W^{-}\to\bar{c}s$ transition. For $P_{T}$, $P_{\bar{D}}$ and $PA_{\bar{B}}$ amplitudes, they are penguin and penguin-annihilation amplitudes, where the subscripts denote the particles receiving the $s$ quark from the $b\to s\bar{c}c$ transition. The effective Hamiltonians governing $\Delta S=0$ transitions can be obtained from the above equations with $H$ replaced with $H^{\prime}$ given in Eq. (53), and with $T_{\bar{D}},C_{\bar{D}T},E_{\bar{D}\bar{B}}$, $P_{T}$, $P_{\bar{D}}$ and $PA_{\bar{B}}$ replaced with $T^{\prime}_{\bar{D}},C^{\prime}_{\bar{D}},E^{\prime}_{\bar{D}\bar{B}}$, $P^{\prime}_{T}$, $P^{\prime}_{\bar{D}}$ and $PA^{\prime}_{\bar{B}}$, respectively. It is understood that these amplitudes are related by CKM matrix elements, giving $T^{\prime}_{\bar{D}}=\frac{V^{*}_{cd}}{V^{*}_{cs}}T_{\bar{D}}$, $C^{\prime}_{\bar{D}T}=\frac{V^{*}_{cd}}{V^{*}_{cs}}C_{\bar{D}T}$, $E^{\prime}_{\bar{D}\bar{B}}=\frac{V^{*}_{cd}}{V^{*}_{cs}}E_{\bar{D}\bar{B}}$, $P^{\prime}_{\bar{D}}=\frac{V^{*}_{td}}{V^{*}_{ts}}P_{\bar{D}}$, $P^{\prime}_{T}=\frac{V^{*}_{td}}{V^{*}_{ts}}P_{T}$ and $PA^{\prime}_{\bar{B}}=\frac{V^{*}_{td}}{V^{*}_{ts}}PA_{\bar{B}}$.
The relative sizes of the tree topological amplitudes are expected to be
$|T_{\bar{D}}|>|C_{\bar{D}T}|>|P_{T}|,|P_{\bar{D}}|,|E_{\bar{D}\bar{B}}|,|PA_{\bar{B}}|.$ (56)
@@
Although the external $W$ diagram is expected to dominate, it is unable to produce a flavor exotic $T$, as the $q\bar{q}$ pair in the diagram, see Fig. 5 (b), is flavor connected. The $\overline{B}_{q}\to\overline{D}T$ decay amplitudes of modes involving $T$ in $\bf 6$} in scenario I or $\overline{\bf 15}$ in scenario II only have contribution from $(C_{\bar{D}T}+P_{T})$, [^2] These two tpological amplitudes can only occur in this combination, see Eqs. (54) and (55). i.e. the diagrams depicted in 5 (a) and (d). Consequently, the amplitudes of these states are highly related.
@@
From the above discussion, we may wrongly have the impression that a $\overline{B}\to\overline{D}T$ decay is similar to a $\overline{B}\to D\overline{T}$ decay. In fact, by comparing Fig. 5 to Fig. 4, one can easily see the differences. For example, in $\overline{B}\to D\overline{T}$ decays, the external $W$ diagram, $T_{\bar{T}}$, has a $\overline{T}$ emitted from $W^{-}$, as shown in Fig. 4 (b), while in $\overline{B}\to\overline{D}T$ decays, the emitted particle is $\overline{D}$ instead, see Fig. 5 (b). Furthermore, there is no counterpart of the internal $W$ diagram $C_{\bar{D}T}$, Fig. 5 (a), in $\overline{B}\to D\overline{T}$ decays, as the $C_{\bar{T}}$ in Fig. 4 (a) is different from $C_{\bar{D}T}$. Explicitly, the former has $\bar{D}T$ sharing the $c\bar{q}$ from $W$, while in the latter, both $c\bar{q}$ go to $\overline{T}$. Consequently, one cannot identify the topological amplitudes in $\overline{B}\to D\overline{T}$ to those in $\overline{B}\to\overline{D}T$ decays.
@@ IV.2.1 $\overline{B}\to D\overline{T}$ decay amplitudes
# Mode $A(\overline{B}_{q}\to D\overline{T}_{\bar{c}\bar{q}[q^{\prime}q^{\prime\prime}]})$ # Mode $A(\overline{B}_{q}\to D\overline{T}_{\bar{c}\bar{q}[q^{\prime}q^{\prime\prime}]})$
$\bar{1}^{*}$ $B^{-}\to D^{+}\overline{T}^{\,--}_{\bar{c}s}$ $(C_{\bar{T}}+P_{\bar{T},1})$ $\bar{2}^{*}$ $B^{-}\to D^{0}\overline{T}^{\,-}_{\bar{c}s}$ $\frac{1}{\sqrt{2}}(C_{\bar{T}}+P_{\bar{T},1})$
$\bar{2}^{*}$ $\overline{B}{}^{0}\to D^{+}\overline{T}^{\,-}_{\bar{c}s}$ $\frac{1}{\sqrt{2}}(C_{\bar{T}}+P_{\bar{T},1})$ $\bar{3}^{*}$ $\overline{B}{}^{0}\to D^{0}\overline{T}^{0}_{\bar{c}s}$ $(C_{\bar{T}}+P_{\bar{T},1})$
$\bar{5}$ $\overline{B}{}_{s}^{0}\to D^{+}\overline{T}^{\,-}_{\bar{c}}$ $\frac{1}{\sqrt{2}}(C_{\bar{T}}+P_{\bar{T},1})$ $\bar{6}$ $\overline{B}{}_{s}^{0}\to D^{0}\overline{T}^{0}_{\bar{c}}$ $\frac{1}{\sqrt{2}}(C_{\bar{T}}+P_{\bar{T},1})$
$\bar{7}$ $\overline{B}{}_{s}^{0}\to D^{+}\overline{T}^{\,\prime\prime-}_{\bar{c}}$ $\frac{1}{\sqrt{2}}((C_{\bar{T}}+P_{\bar{T},1})+2(E_{\bar{T}\bar{B}}+PA_{\bar{B}}))$ $\bar{8}$ $\overline{B}{}_{s}^{0}\to D^{0}\overline{T}^{\,\prime\prime 0}_{\bar{c}}$ $\frac{1}{\sqrt{2}}((C_{\bar{T}}+P_{\bar{T},1})+2(E_{\bar{T}\bar{B}}+PA_{\bar{B}}))$
$\bar{9}$ $B^{-}\to D^{0}\overline{T}^{\,\prime\prime-}_{\bar{c}s}$ $\frac{1}{\sqrt{2}}((C_{\bar{T}}+P_{\bar{T},1})+2(T_{\bar{T}}+P_{\bar{T},2}))$ $\bar{9}$ $\overline{B}{}^{0}\to D^{+}\overline{T}^{\,\prime\prime-}_{\bar{c}s}$ $\frac{1}{\sqrt{2}}((C_{\bar{T}}+P_{\bar{T},1})+2(T_{\bar{T}}+P_{\bar{T},2}))$
$\bar{9}$ $\overline{B}{}_{s}^{0}\to D_{s}^{+}\overline{T}^{\,\prime\prime-}_{\bar{c}s}$ $\sqrt{2}(-T_{\bar{T}}-P_{\bar{T},2}+E_{\bar{T}\bar{B}}+PA_{\bar{B}})$
# Mode $A^{\prime}(\overline{B}_{q}\to D\overline{T}_{\bar{c}\bar{q}[q^{\prime}q^{\prime\prime}]})$ # Mode $A^{\prime}(\overline{B}_{q}\to D\overline{T}_{\bar{c}\bar{q}\{q^{\prime}q^{\prime\prime}\}})$
$\bar{1}^{*}$ $B^{-}\to D_{s}^{+}\overline{T}^{\,--}_{\bar{c}s}$ $(C^{\prime}_{\bar{T}}+P^{\prime}_{\bar{T},1})$ $\bar{2}^{*}$ $\overline{B}{}^{0}\to D_{s}^{+}\overline{T}^{\,-}_{\bar{c}s}$ $\frac{(C^{\prime}_{\bar{T}}+P^{\prime}_{\bar{T},1})}{\sqrt{2}}$
$\bar{4}^{*}$ $\overline{B}{}_{s}^{0}\to D^{0}\overline{T}^{0}_{\bar{c}\bar{s}}$ $(C^{\prime}_{\bar{T}}+P^{\prime}_{\bar{T},1})$
$\bar{5}$ $B^{-}\to D^{0}\overline{T}^{\,-}_{\bar{c}}$ $\frac{1}{\sqrt{2}}(C^{\prime}_{\bar{T}}+P^{\prime}_{\bar{T},1})$ $\bar{5}$ $\overline{B}{}_{s}^{0}\to D_{s}^{+}\overline{T}^{\,-}_{\bar{c}}$ $\frac{1}{\sqrt{2}}(C^{\prime}_{\bar{T}}+P^{\prime}_{\bar{T},1})$
$\bar{6}$ $\overline{B}{}^{0}\to D^{0}\overline{T}^{0}_{\bar{c}}$ $\frac{1}{\sqrt{2}}(C^{\prime}_{\bar{T}}+P^{\prime}_{\bar{T},1})$
$\bar{7}$ $B^{-}\to D^{0}\overline{T}^{\,\prime\prime-}_{\bar{c}}$ $\frac{1}{\sqrt{2}}((C^{\prime}_{\bar{T}}+P^{\prime}_{\bar{T},1})+2(T^{\prime}_{\bar{T}}+P^{\prime}_{\bar{T},2}))$ $\bar{7}$ $\overline{B}{}^{0}\to D^{+}\overline{T}^{\,\prime\prime-}_{\bar{c}}$ $\sqrt{2}((E^{\prime}_{\bar{T}\bar{B}}+PE^{\prime}_{\bar{B}})-(T^{\prime}_{\bar{T}}+P^{\prime}_{\bar{T},2}))$
$\bar{7}$ $\overline{B}{}_{s}^{0}\to D_{s}^{+}\overline{T}^{\,\prime\prime-}_{\bar{c}}$ $\frac{1}{\sqrt{2}}((C^{\prime}_{\bar{T}}+P^{\prime}_{\bar{T},1})+2(T^{\prime}_{\bar{T}}+P^{\prime}_{\bar{T},2}))$ $\bar{8}$ $\overline{B}{}^{0}\to D^{0}\overline{T}^{\,\prime\prime 0}_{\bar{c}}$ $\frac{1}{\sqrt{2}}((C^{\prime}_{\bar{T}}+P^{\prime}_{\bar{T},1})+2(E^{\prime}_{\bar{T}\bar{B}}+PE^{\prime}_{\bar{B}}))$
$\bar{9}$ $\overline{B}{}^{0}\to D_{s}^{+}\overline{T}^{\,\prime\prime-}_{\bar{c}s}$ $\frac{1}{\sqrt{2}}((C^{\prime}_{\bar{T}}+P^{\prime}_{\bar{T},1})+2(E^{\prime}_{\bar{T}\bar{B}}+PA^{\prime}_{\bar{B}}))$